OpenStudy (anonymous):

What is the closest approximation of the solution of the equation 2x – 1 = 3x + 1? (a) -4.42 (b) -5.81 (c) -3.22 (d) 4.93 (e) 3.33

OpenStudy (alexwee123):

2x – 1 = 3x + 1 -x-1=1 -x=2 x=-2 uhh idk about you but i think the answer sets are wrong....

OpenStudy (anonymous):

questioning a problem?

OpenStudy (alexwee123):

lol i guess :/

OpenStudy (anonymous):

It's asking for the closest approximation...you got x= -2...which is closest to -2?

OpenStudy (anonymous):

That is unless the problem is written incorrectly :)